OpenStudy (anonymous):

4x-3/5=2/5 A.1/5 B.5/5 C.1/2 D.1/4 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

HanAkoSolo (jamierox4ev3r):

4x-3/5=2/5 +3/5 +3/5 ----------- 4x=1 /4 /4 x=1/4 Correct answer is D.